# Cooperative Payload Estimation by a Team of Mocobots

### v1

@@ Abstract
Consider the following scenario: a human guides multiple mobile manipulators to grasp a common payload. For subsequent high-performance autonomous manipulation of the payload by the mobile manipulator team, or for collaborative manipulation with the human, the robots should be able to discover where the other robots are attached to the payload, as well as the payload’s mass and inertial properties. In this paper, we describe a method for the robots to autonomously discover this information. The robots cooperatively manipulate the payload, and the twist, twist derivative, and wrench data at their grasp frames are used to estimate the transformation matrices between the grasp frames, the location of the payload’s center of mass, and the payload’s inertia matrix. The method is validated experimentally with a team of three mobile cobots, or mocobots.
@@ I Introduction
Consider a scenario in manufacturing, logistics, or construction where a large, substantially rigid payload must be manipulated in all six degrees of freedom (dof), perhaps for an assembly or loading task. Multiple distributed contacts with the payload are required, to respect the workspace and wrench limits of any single manipulator and to minimize stress concentrations for heavy or fragile payloads. If the manipulation task is not one that is easily automated (e.g., it is not a repetitive task performed in a structured environment), then one or more human operators can physically collaborate with a team of mobile cobots, or mocobots [1]. The partnership combines the mocobots’ physical strength with human perception and adaptability (Figure 1).
First, a human guides the mobile manipulators to grasp locations on the common payload. Once the grasps are established, to provide optimal model-based assistance to the human, or for high-performance autonomous manipulation, the robots should be able to discover where the other robots are attached to the payload, as well as the payload’s mass and inertial properties.
In previous work [1], much of this information was provided to the mocobots in advance of manipulation. In this paper, we describe a method for the robots to discover this information. The robots cooperatively manipulate the payload, and the twist, twist derivative, and wrench data at their grasp frames are used to estimate the transformation matrices between the grasp frames, the location of the payload’s center of mass, and the payload’s inertia matrix. The method is validated experimentally with a team of three mocobots.
I-A Related Work
A “mocobot” is a mobile variant of the “cobot,” robots designed for physical collaboration with humans, originally introduced in [2]. Mocobots enhance cobots with mobility for diverse tasks. The Omnid mocobot, designed for human-multirobot collaborative manipulation, is described in [1].
The work in this paper extends previous work by providing a method for discovering payload kinematic and inertial properties, building on previous work in cooperative robot manipulation and payload estimation.
I-A1 Cooperative Robot Manipulation
Cooperative robot manipulation involves multiple robots working together to manipulate objects, with applications in manufacturing, construction, and hazardous environments [3, 4, 5, 6]. Various control architectures ranging from centralized to distributed systems, including those with no explicit communication, have been utilized for manipulation tasks [7, 8, 9, 10].
@@ I-A1 Cooperative Robot Manipulation
Approaches like haptic feedback for leader-follower formations [11] and force amplification strategies [12] have been explored. Recent advancements include machine learning techniques to enhance human-robot interaction and adaptability [13, 14]. Some systems use distributed adaptive control of omnidirectional mobile bases equipped with robot arms to autonomously transport payloads [15, 16]. Other configurations include customized robots for collaborative rigid-body manipulation [17] and strategies for manipulating deformable payloads [18].
I-A2 Payload Estimation
Knowledge of payload properties, particularly the center of mass (CoM) and inertia matrix, is crucial for the dynamic performance and stability of robotic systems. Early research on payload inertia matrix estimation involved suspending payloads and measuring oscillation periods, which provided reliable results for vehicles and aircraft [19, 20, 21, 22, 23]. However, these techniques are impractical for robotic applications, especially in dynamic and unstructured environments.
In the field of robotics, force-torque sensors, combined with Newton-Euler equations, have become a standard tool for the estimation of inertial properties by a single robot [24]. For the case of manipulation by multiple mobile manipulators, decentralized approaches have been developed for estimating properties of the common payload [25, 26]. In [25], the robots cooperatively estimate the centroid of the payload, but the motion is limited to the plane and mass and inertial parameters are not considered. The estimation goals of [26] are closely related to those of this paper, except the algorithms described in that work do not consider the orientations of the robots’ grasp frames and have only been tested in simulation.
I-B Contributions
This paper presents a methodology for using a group of mobile robots to discover properties of a rigid-body payload, including the robots’ unknown relative grasp frames on the payload and the payload’s mass, center of mass, and inertia matrix. The approach relies on post-processing experimental data collected without the need for expensive and potentially fragile force-torque sensors. The method is validated experimentally on a team of three custom mocobots.
@@ II Cooperative Rigid Payload Estimation Problem Formulation
The cooperative rigid payload estimation problem can be formulated as follows: given $\mathcal{D}_{**}$, determine the payload’s mass ${\mathfrak{m}}$; the configuration of a frame $\{\operatorname{c}\}$ at the payload’s center of mass relative to each robot’s grasp frame ($T_{1c},\ldots,T_{Nc}$) such that the frame $\{\operatorname{c}\}$ is aligned with principal axes of inertia of the payload; and the $3\times 3$ positive-definite inertia matrix $\mathcal{I}$ of the payload in the frame $\{\operatorname{c}\}$. This information also implies the configuration of each robot’s grasp relative to each other,
@@
III Estimation of Payload Properties Using the Omnid Mocobots
III-A Omnid Mocobots and Sequential Estimation
We perform payload estimation using the Omnid mocobots (Figure 2), which are designed for human-robot collaborative manipulation [1]. Each Omnid mocobot consists of a mecanum-wheel mobile base, a 3-dof Delta parallel mechanism driven by series-elastic actuators (SEAs), and a 3-dof passive gimbal wrist. In this paper, the mobile bases are stationary, and all manipulation is performed by the Delta-plus-gimbal manipulators.
The Delta mechanism and gimbal wrist of each Omnid is equipped with encoders, allowing Omnid $i$ to measure its wrist configuration $T_{i_{0}i}$. Since the payload is rigidly attached to the gimbal, the wrist configuration and the grasp configuration are equivalent. The twist ${\mathcal{V}}_{i}$ and its derivative $\dot{{\mathcal{V}}}_{i}$ are calculated by filtering encoder data. Wrenches at the wrist take the form ${\mathcal{F}}_{i}=(m_{i},f_{i})=(0,f_{i})$, where the linear force $f_{i}$ is calculated based on torques measured at the SEA joints and the moment $m_{i}$ is zero due to the passive gimbal wrist. Data is collected at a fixed sampling rate of 100 Hz.
Refer to caption (a)
Figure 2: The structure of the Omnid mocobot.
In principle, all data $\mathcal{D}_{**}$ could be used in a single optimization to simultaneously calculate the configurations of the robot grasps relative to each other, the mass and location of the center of mass of the payload, and the inertia of the payload. In this paper, we adopt a sequenced approach, where we first estimate the grasp kinematics using only twist measurements; then estimate the mass and center of mass using the results of the kinematics estimates and wrench measurements when the payload is held stationary; and finally estimate the inertial properties using the results of the previous estimates as well as twist, acceleration, and wrench data during manipulation. This approach (a) allows us to use simple least-squares estimation and (b) requires only the data needed for the particular parameters being estimated—for example, wrench and acceleration data is not needed to estimate the robots’ relative grasp frame configurations.
@@ III-B Grasp Kinematics
When robots $i$ and $j$ grasp a common rigid body, the twists ${\mathcal{V}}_{i},{\mathcal{V}}_{j}$ are related by ${\mathcal{V}}_{i}=[\text{Ad}_{T_{ij}}]{\mathcal{V}}_{j}$, where $[\text{Ad}_{T_{ij}}]\in^{6\times 6}$ is the adjoint representation of the transformation matrix $T_{ij}$ describing the configuration of the frame $\{j\}$ relative to $\{i\}$ [27]. We write the complete set of twist measurements as ${\mathcal{V}}_{i*},{\mathcal{V}}_{j*}\in\mathbb{R}^{6\times P}$, i.e., each individual twist measurement ${\mathcal{V}}_{i}$ forms a column of the matrix ${\mathcal{V}}_{i*}$. Then
@@ III-B1 Rotation Matrix
Let $\omega_{i*},\omega_{j*}\in\mathbb{R}^{3\times P}$ be the angular velocities of the twists measured at each grasp. The relationship between the two sets is
@@ III-B2 Position Vector
Let $v_{i*},v_{j*}\in\mathbb{R}^{3\times P}$ be the linear velocities of the twists measured at each grasp. By Equation (2), the linear velocities satisfy
@@ III-C Mass and Center of Mass (CoM)
The mass and center of mass are estimated using the grasp kinematics solution and a set of static wrench measurements. A common reference frame $\{\operatorname{s}\}\in\{\{1\},\ldots,\{N\}\}$ is selected for the estimation process. We also define a frame $\{\operatorname{s}_{0}\}$ coincident with $\{\operatorname{s}\}$ but oriented such that its $\hat{z}$-axis is opposite the gravity vector ${\mathfrak{g}}\in^{3}$. The center of mass of the payload is located at the origin of a to-be-estimated frame $\{\operatorname{c}\}$.
The force and moment static equilibrium conditions are
@@
$[p_{sc}]({R_{ss_{0}}}{\mathfrak{m}}{\mathfrak{g}})$ $=-\sum\limits_{i=1}^{N}[p_{si}](R_{si}f_{i}),$ (10)
@@
To ensure static equilibrium, the payload must be supported by three or more robots, as two Omnids cannot resist a moment about an axis through the two grasp frames. Under this condition, the mass ${\mathfrak{m}}$ can be estimated using Equation (9) and one or more static measurements by the robots, while $p_{sc}$ can be estimated using Equation (10) and two or more static measurements, provided the orientations of the payload during measurement differ by rotation about an axis not aligned with the gravity vector ${\mathfrak{g}}$.
@@
$A_{q}=[{R_{ss_{0}}^{q}}\hat{{\mathfrak{m}}}{\mathfrak{g}}]\in^{3\times 3},\quad b_{q}=\sum_{i=1}^{N}[p_{si}](R_{si}^{q}f_{i}^{q})\in^{3\times 1},$ (14)
@@ III-D Inertia Matrix
We define a frame $\{\operatorname{b}\}$ at the center of mass $\hat{p}_{sc}$ and aligned with $\{\operatorname{s}\}$. After estimating the inertia matrix in $\{\operatorname{b}\}$, we identify the principal axes of inertia in $\{\operatorname{b}\}$, and then define a final center-of-mass frame $\{\operatorname{c}\}$ with axes aligned with the principal axes of inertia.
@@
where $m_{b}$ is the moment acting on the body, $\mathcal{I}_{b}\in\mathbb{R}^{3\times 3}$ is the positive-definite inertia matrix of the payload, $\omega_{b}$ is the angular velocity, and $\alpha_{b}$ is the angular acceleration, all expressed in the frame $\{\operatorname{b}\}$. For $N$ robots supporting the payload via linear forces $f_{i}$ at the gimbals, the moment can be expressed as
$m_{b}=\sum_{i=1}^{N}[p_{bi}](R_{bi}f_{i}).$ (18)
@@
IV Experimental Implementation
@@ IV-A Experimental Configuration
The payload used in the experimental validation is a frame made from ready-tube aluminum extrusion. The payload is attached to the Omnid gimbal wrists, and the attachments remain unchanged during experiments. The overall setup is shown in Figure 3. The Omnid grasp frames are denoted $\{1\}$, $\{2\}$, and $\{3\}$. Frame $\{1\}$ is selected as the reference frame $\{\operatorname{s}\}$.
Each 6-dof manipulator consists of a 3-dof Delta mechanism and a 3-dof passive gimbal joint (Figure 4). When the payload is rigidly attached, the top of the gimbal wrist is considered part of the payload. Accordingly, the grasp frame $\{i\}$ is positioned at the center of the gimbal wrist.
Figure 3: Frame definition of the robot-payload system for experiments. Frame $\{1\}$ is chosen as the reference frame $\{\operatorname{s}\}$. The axes of the frames are visually differentiated by colors with red corresponding to the $x$-axis, green to the $y$-axis, and blue to the $z$-axis.
@@
Rotation Matrix in Axis Angle (deg) $R_{12}$ $R_{23}$ $R_{31}$
$(0,0,0)$ $(0,0,90)$ $(0,0,-90)$
Position Vector (m) $p_{12}$ $p_{23}$ $p_{31}$
$(0.647,0.533,0)$ $(0.457,-0.838,0)$ $(0.305,1.105,0)$
Mass (kg), Center of Mass (m), and Principal Axes (deg) ${\mathfrak{m}}$ $p_{1c}$ $R_{1c}$
$11.412$ $(0.553,0.172,0.072)$ $(0.001,0.010,-95.016)$
@@
$2.258$ $3.448$ $5.698$
TABLE I: Ground truth values of the experimental configuration.

### v3

@@ Abstract
For high-performance autonomous manipulation of a payload by a mobile manipulator team, or for collaborative manipulation with the human, robots should be able to discover where other robots are attached to the payload, as well as the payload’s mass and inertial properties. In this paper, we describe a method for the robots to autonomously discover this information. The robots cooperatively manipulate the payload, and the twist, twist derivative, and wrench data at their grasp frames are used to estimate the transformation matrices between the grasp frames, the location of the payload’s center of mass, and the payload’s inertia matrix. The method is validated experimentally with a team of three mobile cobots, or mocobots.
@@ I Introduction
Consider a scenario in manufacturing, logistics, or construction where a large, substantially rigid payload must be manipulated in all six degrees of freedom (DoF), perhaps for an assembly or loading task. Multiple distributed contacts with the payload are required to respect the wrench limits of any single manipulator and to minimize stress concentrations for heavy or fragile payloads. If the manipulation task is not one that is easily automated (e.g., it is not a repetitive task performed in a structured environment), then one or more human operators can physically collaborate with a team of mobile cobots, or mocobots [1]. The partnership combines the mocobots’ physical strength with human perception and adaptability (Figure 1).
First, a human guides the mobile manipulators to grasp the common payload. Once the grasps are established, to provide optimal model-based assistance to the human, or for high-performance autonomous manipulation, the robots should be able to discover where the other robots are attached to the payload, as well as the payload’s mass and inertial properties.
In this paper, we describe a method for the robots to autonomously discover this information. The robots cooperatively manipulate the payload, and the twist, twist derivative, and wrench data at their grasp frames are used to estimate the transformation matrices between the grasp frames, the location of the payload’s center of mass, and the payload’s inertia matrix. The method is validated experimentally with a team of three mocobots.
@@
I-A Related Work
A “mocobot” is a mobile variant of a “cobot,” robots designed for physical collaboration with humans, originally introduced in [2]. Mocobots enhance cobots with mobility for diverse tasks. An example is the Omnid mocobot, designed for human-multirobot collaborative manipulation [1].
Cooperative robot manipulation involves multiple robots working together to manipulate objects, with a variety of applications in manufacturing, construction, and hazardous environments [3, 4, 5, 6]. Accurate knowledge of a payload’s inertial properties enables planning for dynamic manipulation and high-performance control [7, 8, 9]. Various cooperative control architectures, including centralized and decentralized controllers, have been proposed for manipulation and transportation under the assumption of known inertial properties [10, 11, 12]. This assumption motivates estimation of such properties for unknown payloads.
A recent survey [7] provides a comprehensive review of estimation of payload inertial properties, divided into three main categories: purely visual methods, exploratory methods, and methods where the payload is rigidly attached to or grasped by a robot manipulator. The purely visual method relies only on static images or video of the object. For example, the inertial parameters of an object can be estimated visually from the size and shape of the object under assumptions on its density properties [13]. Most visual methods require prior knowledge or assumptions, however, such as knowledge of the total mass of the object or an assumption of uniform density. Exploratory methods include pushing, poking, or tilting an object (e.g., [14, 15]), but motions are often limited to a plane and only a subset of inertial properties can be estimated. When a payload is firmly attached to a robot manipulator, Newton-Euler equations can be used to estimate payload inertial properties from robot motion and wrist force-torque data [16]. The basic idea can be extended to grasps by multiple robots, as in [17, 18] for a planar payload and [19] for a 3D payload. The approach and goals of [19] are closely related to those of this paper, except [19] does not consider the orientations of the robots’ grasp frames and the approach was tested only in simulation.
I-B Contribution
This paper presents a methodology enabling a group of mobile manipulators to estimate properties of an unknown rigid-body payload. These properties include the payload’s mass, center of mass, and inertia matrix, as well as the transformation matrices between the robots’ grasp frames, which are initially unknown. We have performed extensive experimental validation with mocobots and their force-controlled manipulators, which are well suited to cooperative manipulation and do not have expensive force-torque sensors at their wrists. Results from these experiments highlight the method’s robustness and practicality.
@@ II Cooperative Rigid Payload Estimation Problem Formulation
The cooperative rigid payload estimation problem can be formulated as follows: given $\mathcal{D}_{**}$, determine the payload’s mass ${\mathfrak{m}}$; the configuration of a frame $\{\operatorname{c}\}$ at the payload’s center of mass relative to each robot’s grasp frame ($T_{1c},\ldots,T_{Nc}$) such that the frame $\{\operatorname{c}\}$ is aligned with principal axes of inertia of the payload; and the $\times 3$ 3 positive-definite inertia matrix $\mathcal{I}$ of the payload in the frame $\{\operatorname{c}\}$. This information also implies the configuration of each robot’s grasp relative to the others,
@@
III Estimation of Payload Properties
III-A Sequential Estimation
In principle, all data $\mathcal{D}_{**}$ could be used in a single optimization to simultaneously calculate the configurations of the robot grasps relative to each other, the mass and location of the center of mass of the payload, and the inertia of the payload. In this paper, we adopt a sequenced approach, where we first estimate the grasp kinematics using only twist measurements; then estimate the mass and center of mass using the results of the kinematics estimates and wrench measurements when the payload is held stationary; and finally estimate the inertial properties using the results of the previous estimates as well as twist, acceleration, and wrench data during manipulation. This approach (a) allows us to use simple least-squares estimation and (b) requires only the data needed for the particular parameters being estimated; for example, wrench and acceleration data are not needed to estimate the robots’ relative grasp frame configurations.
@@ III-B Grasp Kinematics
When robots $i$ and $j$ grasp a common rigid body, the twists ${\mathcal{V}}_{i}$ and ${\mathcal{V}}_{j}$ are related by ${\mathcal{V}}_{i}=[\text{Ad}_{T_{ij}}]{\mathcal{V}}_{j}$, where $[\text{Ad}_{T_{ij}}]\in^{6\times 6}$ is the adjoint representation of the transformation matrix $T_{ij}$ describing the configuration of the frame $\{j\}$ relative to $\{i\}$ [20]. Frames are illustrated in Figure 2. We write the complete set of twist measurements as ${\mathcal{V}}_{i*},{\mathcal{V}}_{j*}\in\mathbb{R}^{6\times Q}$, i.e., each individual twist measurement ${\mathcal{V}}_{i}$ forms a column of the matrix ${\mathcal{V}}_{i*}$. Then
@@ III-B1 Rotation Matrix
Let $\omega_{i*},\omega_{j*}\in\mathbb{R}^{3\times Q}$ be the angular velocities of the twists measured at each grasp. The relationship between the two sets is
@@ III-B2 Position Vector
Let $v_{i*},v_{j*}\in\mathbb{R}^{3\times Q}$ be the linear velocities of the twists measured at each grasp. By Equation (2), the linear velocities satisfy
@@ III-C Mass and Center of Mass (CoM)
[…]
With local wrenches ${\mathcal{F}}_{i}=(m_{i},f_{i})$ from each robot, the force and moment static equilibrium conditions are
@@
$[p_{sc}]({R_{ss_{0}}}{\mathfrak{m}}{\mathfrak{g}})=-\sum\limits_{i=1}^{N}\Big{(}[p_{si}](R_{si}f_{i})+R_{si}m_{i}\Big{)},$ (10)
@@
The mass ${\mathfrak{m}}$ can be estimated using Equation (9) and one or more static measurements by the robots, while $p_{sc}$ can be estimated using Equation (10) and two or more static measurements, provided the orientations of the payload during measurement differ by rotation about an axis not aligned with the gravity vector ${\mathfrak{g}}$.
@@
$A_{q}=[{R_{ss_{0}}^{q}}\hat{{\mathfrak{m}}}{\mathfrak{g}}]\in^{3\times 3},$ (14)
$b_{q}=\sum_{i=1}^{N}\Big{(}[p_{si}](R_{si}^{q}f_{i}^{q})+R_{si}^{q}m_{i}^{q}\Big{)}\in\mathbb{R}^{3\times 1},$ (15)
@@ III-D Inertia Matrix
We define a frame $\{\operatorname{b}\}$, shown in Figure 2, at the center of mass $\hat{p}_{sc}$ and aligned with $\{\operatorname{s}\}$. After estimating the inertia matrix in $\{\operatorname{b}\}$, we identify the principal axes of inertia in $\{\operatorname{b}\}$, and then define a final center-of-mass frame $\{\operatorname{c}\}$ with axes aligned with the principal axes of inertia.
@@
where $m_{b}$ is the moment acting on the body, $\mathcal{I}_{b}\in\mathbb{R}^{3\times 3}$ is the positive-definite inertia matrix of the payload, $\omega_{b}$ is the angular velocity, and $\alpha_{b}$ is the angular acceleration, all expressed in the frame $\{\operatorname{b}\}$. For $N$ robots supporting the payload with local wrenches ${\mathcal{F}}_{i}=(m_{i},f_{i})$, the moment can be expressed as
$m_{b}=\sum_{i=1}^{N}\Big{(}[p_{bi}](R_{bi}f_{i})+R_{bi}m_{i}\Big{)}.$ (19)
@@
IV Experimental Implementation with the Omnid Mocobots
Refer to caption (a)
Figure 3: The structure of the Omnid mocobot.
We performed payload estimation using the Omnid mocobots (Figure 3), which are designed for human-robot collaborative manipulation [1]. Each Omnid mocobot consists of a mecanum-wheel mobile base, a 3-DoF Delta parallel mechanism driven by series-elastic actuators (SEAs), and a 3-DoF passive gimbal wrist. In this paper, the mobile bases are stationary, and all manipulation is performed by the Delta-plus-gimbal manipulators.
The Delta mechanism and gimbal wrist of each Omnid are equipped with encoders, allowing Omnid $i$ to measure its wrist configuration $T_{i_{0}i}$. Since the payload is rigidly attached to the gimbal, the wrist configuration and the grasp configuration are equivalent. The twist ${\mathcal{V}}_{i}$ and its derivative $\dot{{\mathcal{V}}}_{i}$ are calculated by filtering encoder data. Wrenches at the wrist take the form ${\mathcal{F}}_{i}=(m_{i},f_{i})=(0,f_{i})$, where the linear force $f_{i}$ is calculated based on torques measured at the SEA joints and the moment $m_{i}$ is zero due to the passive gimbal wrist. Data is collected at a fixed sampling rate of 100 Hz.
[…]
@@ IV-A Experimental Configuration
To experimentally validate the methodology, we used four different payloads: two structures constructed from aluminum extrusion, each with two different sets of grasp locations (Figure 4).
[…]
@@
Figure 5: Experimental setup with payload configuration (d).
Rotation Matrix in Axis Angle (deg)
Config. $R_{12}$ $R_{23}$ $R_{31}$
(a) (0, 0, 0) (0, 0, --90) (0, 0, 90)
(b) (0, 0, 45) (0, 0, 90) (0, 0, --135)
(c) (0, 0, --90) (0, 0, 45) (0, 0, 45)
(d) (0, 0, --135) (0, 0, 45) (0, 0, 90)
Position Vector (m)
Config. $p_{12}$ $p_{23}$ $p_{31}$
(a) (0.647, 0.533, 0) (0.457, --0.838, 0) (--0.305, --1.105, 0)
(b) (--0.686, 0.572, 0) (0.916, --0.647, 0) (--0.242, 0.835, 0)
(c) (--0.115, --1.143, 0) (--0.229, 0.800, 0) (--1.132, 0.162, 0)
(d) (--0.382, 0.800, --0.038) (--0.485, 0.862, 0.038) (0.533, --0.571, 0)
Mass (kg), Center of Mass (m), and Principal Axes (deg)
Config. ${\mathfrak{m}}$ $p_{1c}$ $R_{1c}$
(a) 11.672 (0.537, 0.156, 0.072) (0.043, 0.085, -84.787)
(b) 11.672 (0.034, 0.535, 0.072) (0.115, --0.115, 2.465)
(c) 10.478 (0.547, --0.670, 0.057) (--1.034, 0.355, --24.643)
(d) 10.478 (0.089, 0.593, 0.055) (--1.662, --0.702, 68.937)
@@
Config. ${\mathcal{I}}_{xx}$ ${\mathcal{I}}_{yy}$ ${\mathcal{I}}_{zz}$
(a) 2.318 3.215 5.524
(b) 2.153 3.390 5.535
(c) 1.824 2.438 4.208
(d) 1.711 2.122 3.772
TABLE I: Ground truth values for the four experimental configurations.